# FinSphere: A Conversational Stock Analysis Agent Equipped with Quantitative Tools based on Real-Time Database

### v1

@@ 5.2 Ablation Study
To investigate the impact of training data scale on FinSphere’s performance, we conduct an ablation study using different proportions of the Stocksis dataset. We fine-tune Qwen2-72B using 20%, 50%, 80%, and 100% of the 5,000 data pairs while maintaining FinSphere’s framework. The detailed evaluation results are provided in Appendix B.
The results demonstrate a clear positive correlation between training data scale and model performance, with overall scores increasing from 58.90 (20%) to 70.88 (100%). Notably, the performance improvement shows a non-linear pattern, with larger incremental gains observed at higher data volumes (3.78 points from 20% to 50%, 3.99 points from 50% to 80%, and 4.21 points from 80% to 100%). These findings underscore the importance of comprehensive training data in achieving optimal performance, while also demonstrating the robust scalability of our framework, as FinSphere maintains satisfactory performance levels even with reduced training data.
@@ 6 Conclusion
This paper introduces FinSphere, an innovative stock analysis agent that addresses critical gaps in the capabilities of LLMs for stock analysis. By integrating real-time financial databases, quantitative tools, and an instruction-tuned LLM, FinSphere demonstrates superior performance in generating comprehensive stock analyses. The development and release of Stocksis, a high-quality dataset for enhancing LLMs’ stock analysis capabilities, and AnalyScore, a systematic evaluation framework, provide valuable resources for advancing research in AI-powered financial analysis. Our experimental results indicate that FinSphere consistently outperforms general-purpose, domain-specific LLMs and Agent systems across multiple evaluation dimensions, highlighting the effectiveness of our integrated approach. This work represents a significant advancement toward democratizing access to professional-grade financial analysis tools while maintaining analytical rigor and practical utility.
@@ Appendix A AnalyScore Details
AnalyScore introduced in Section 3.2 is fully presented in this section, hoping to promote related research in academia and industry. Please refer to Table 4 for the priori eligibility check and Table 5 for the specific composition of AnalyScore.
Appendix B Ablation Results
We put the result table of the ablation experiment in the main text here, for details, please see Table 3.
Appendix C More Analysis and Visualization
In our analysis in Section 5.1, we have demonstrated that agent-based systems generally outperform single LLM in stock analysis. To further examine the relative strengths of FinSphere, we have conducted a detailed dimensional comparison between FinSphere and two prominent agent-based systems (FinRobot and FinMem). The comparative visualization in Figure 2 reveals the nuanced performance differences across four key dimensions: Conclusion, Content, Expression, and Data capabilities.
From the visualization, we can observe FinSphere’s superior performance across all dimensions. In the Conclusion dimension, the three models show relatively close scores, with FinSphere slightly leading at 9.95, followed by FinMem at 9.90 and FinRobot at 9.10, indicating their reliable capability in deriving investment conclusions. The Content dimension reveals more significant differences, where FinSphere notably outperforms with 27.16 points compared to FinMem (25.95) and FinRobot (24.05), demonstrating FinSphere’s distinct advantage in analysis depth and content richness. The Expression dimension shows the most pronounced gap, with FinSphere scoring 14.87, significantly higher than FinMem (12.85) and FinRobot (11.55), highlighting FinSphere’s excellence in logical articulation and clarity of analysis. In the Data dimension, FinSphere and FinMem perform similarly (18.90 and 18.85 respectively), both substantially outperforming FinRobot (16.35), indicating superior data processing and utilization capabilities in the former two models.
@@ Appendix C More Analysis and Visualization
Appendix D Testing Prompt
@@ D.1 Response Standards
  • Technical Analysis: Technical patterns, indicators with specific values. Note specialized indicators if present (e.g., "AI Top/Bottom", "Bull Institution signals").
@@ D.2 Writing Guidelines
  • Avoid phrases: "according to", "information shows", "recent performance", "current situation", "comprehensive analysis", "buy", "sell", "hold", "clear position", "build position", "reduce position", "increase position"
  • Minimize transition words: "but", "however", "then", "finally"
@@
Conclusion Structure Logical Consistency Factual Support
Clear conclusions addressing: $\bullet$ Time-based perspectives $\bullet$ Multi-dimensional analysis Consistency between: $\bullet$ Conclusions and dimensions $\bullet$ Dimensions and data Analysis must include: $\bullet$ Exact and quantifiable data $\bullet$ Specific metrics or systems
Data Timeliness Analytical Dimensions Neutral Language
Requirements: $\bullet$ Technical: Past month $\bullet$ Capital flows: Past 15 days $\bullet$ Financials: Past 3 years $\bullet$ News: Past 15 days At least include one of: $\bullet$ Short-term analysis $\bullet$ Medium/Long-term analysis $\bullet$ Supporting data categories Requirements: $\bullet$ Avoid trading recommendations $\bullet$ Focus on objective analysis $\bullet$ Exclude trading keywords
Table 4: Priori eligibility check of AnalyScore
@@
Data (Total Score: 20) Extensive data (>3 dimensions) 20
@@
Appendix E LLMs’ Response to users’ query
@@ E.1 FinSphere
    • Comprehensive data dimensions (>3 dimensions)
@@ E.2 GPT-4o with few-shot
GPT-4o with few-shot’s analysis begins with vague conclusions: "worth moderate attention in short term, should pay appropriate attention in medium to long term" lacks a clear stance; while data citations are complete, the analysis lacks depth; technical analysis is oversimplified and fails to fully interpret indicator meanings; the conclusion section lacks sufficient reasoning for investment recommendations. Specific scores are as follows:
@@
E.3 QWen2-72B with few-shot
@@
Appendix F A complete example of Stocksis
@@
Table 6: A complete example of Stocksis. Same sample as in Table 1. The average cost of comprehensive analysis written by experts is \$10 per query, with detailed cost breakdowns available in Appendix G.
Appendix G Dataset Collection and Evaluation Costs
To further quantify the value of Stocksis and AnalyScore, we disclose the expert-curated Stocksis dataset containing 5,000 entries and the human evaluation costs for assessing model outputs using AnalyScore. Excluding the costs of quantitative tool calls, the expert compilation of comprehensive analyses costs approximately \$75,000 (averaging \$15 per entry). The expert evaluation of 100 outputs from each of the models using AnalyScore criteria cost approximately \$11,000 (averaging \$10 per evaluation). These figures do not include the additional costs associated with expert development of the AnalyScore evaluation framework.
@@ Appendix G Dataset Collection and Evaluation Costs
Appendix H Product Release Information
As FinSphere—a powerful stock analysis agent developed by a stock investment advisory company—we currently have a fully functional product demo and plan to make it freely available to the public in December 2024. Due to double-blind review requirements, we regrettably cannot showcase this promising product to the reviewers at this stage. However, we look forward to including access information for this free public tool in the final version of our paper.
Appendix I Detailed testing quiries
Here we disclose 100 queries used for testing and experts’ scores on FinSphere. For details, please check the Table 7, 8 and 9.

### v2

@@ 4.3 Ablation Study
Figure 2: Changes in scores of each sub-item as the number of Stocksis used for fine-tuning changes.
To investigate the impact of training data scale on FinSphere’s performance, we conduct an ablation study using different proportions of the Stocksis dataset. We fine-tune Qwen2-72B using 20%, 50%, 80%, and 100% of the 5,000 data pairs while maintaining FinSphere’s framework. The detailed evaluation results are provided in Table 2, which demonstrates a clear positive correlation between training data scale and model performance, with overall scores increasing from 58.90 (20%) to 70.88 (100%). The performance improvement shows a non-linear pattern, with larger incremental gains observed at higher data volumes. These findings underscore the importance of training data in achieving optimal performance, while also demonstrating the robustness of our framework, as FinSphere maintains satisfactory performances even with reduced training data scale.
5 Related Works
LLMs have shown strong potential in financial tasks such as stock prediction, market analysis, and portfolio management (Zhao et al., 2024; Li et al., 2023b; Ni et al., 2024; Bhat and Jain, 2024; Wu, 2024; Kim et al., 2024). Domain-specific models like InvestLM (Yang et al., 2023b) and GPT-InvestAR (Gupta, 2023) further highlight the benefits of financial instruction tuning. LLMs have also been applied to financial anomaly detection (Park, 2024) and financial statement understanding (Kim et al., 2024).
Financial Datasets and Evaluation
Existing datasets such as FinQA (Chen et al., 2021), TAT-QA (Zhu et al., 2021), and FLARE (Xie et al., 2023) focus on question answering or numerical reasoning, but lack support for comprehensive, real-time stock analysis. Others, including FinTextQA (Chen et al., 2024), CFBenchmark (Lei et al., 2023), and FinanceBench (Islam et al., 2023), provide broader coverage but miss expert-annotated stock reports. Evaluation of financial text has typically relied on generic metrics (e.g., BLEU (Papineni et al., 2002), ROUGE (Rouge, 2004)), while recent efforts like FinEval (Zhang et al., 2023b) offer domain-adapted metrics but lack systematic expert-grounded scoring.
Instruction Tuning and Tool-Augmented Agents
Recent advances integrate financial LLMs with domain-specific tools, such as FinGPT (Yang et al., 2023a), XBRL-Agent (Han et al., 2024), and FinOps (Li et al., 2023a), improving task relevance and interactivity. However, most rely on static historical data and lack real-time adaptability. In contrast, our proposed FinSphere (Section 3) combines an instruction-tuned LLM with real-time databases and quantitative tools for dynamic and actionable stock analysis.
@@ 6 Conclusion
This paper introduces FinSphere, an innovative stock analysis agent that addresses critical gaps in the capabilities of LLMs for stock analysis. By integrating real-time financial databases, quantitative tools, and an instruction-tuned LLM, FinSphere demonstrates superior performance in generating comprehensive stock analyses. The development and release of Stocksis, a high-quality dataset for enhancing LLMs’ stock analysis capabilities, and AnalyScore, a systematic evaluation framework, provide valuable resources for advancing research in AI-powered financial analysis. Our experimental results indicate that FinSphere consistently outperforms general-purpose, domain-specific LLMs and Agent systems across multiple evaluation dimensions, highlighting the effectiveness of our integrated approach.
@@ Appendix A AnalyScore Details
AnalyScore, introduced in Section 2.1, is fully presented in this section, hoping to promote related research in academia and industry. Please refer to Table 4 for the specific composition of AnalyScore.
Appendix B Dimensional Analysis and Visualization
To further investigate the comparative strengths of FinSphere, we conduct a detailed analysis of its performance relative to two other leading agent-based systems, FinRobot and FinMem. The comparative visualization in Figure 3 highlights performance differences across four critical dimensions: Conclusion, Content, Expression, and Data capabilities.
@@ Appendix B Dimensional Analysis and Visualization
From the visualization, FinSphere exhibits the highest scores across all dimensions. In the Conclusion category, the three models perform relatively closely, with FinSphere slightly leading at 9.95, followed by FinMem at 9.90 and FinRobot at 9.10, demonstrating their robust ability to derive investment insights. However, in the Content dimension, FinSphere shows a clear advantage, scoring 27.16, significantly surpassing FinMem (25.95) and FinRobot (24.05), reflecting its greater analytical depth and content richness.
The most pronounced gap is observed in the Expression dimension, where FinSphere achieves 14.87, noticeably higher than FinMem (12.85) and FinRobot (11.55). This highlights FinSphere’s superior logical articulation in financial reporting. In terms of Data utilization, FinSphere (18.90) and FinMem (18.85) exhibit comparable performance, both substantially outperforming FinRobot (16.35), reinforcing their accurate understanding and grasp of data in financial analysis
On the other hand, one notable limitation of general-purpose LLMs is their heavy dependence on extensive in-context examples to generate accurate financial analyses. This results in a substantial increase in input token consumption, leading to higher operational costs for models such as GPT-4o, and restricting the usability of small context-window LLMs. In contrast, FinSphere’s instruction-tuned architecture eliminates the need for verbose prompts, allowing it to generate high-quality outputs with significantly fewer input tokens.
The combined findings underscore FinSphere’s state-of-the-art capabilities in stock analysis, driven by its robust data processing and structured analytical reasoning. These results further validate its advantage over both standalone LLMs and agent-based financial analysis systems. Additionally, FinSphere has been released to the public with free access in December 2024, with detailed release information provided in Appendix G.
Appendix C Testing Prompt
@@ C.1 Response Standards
  • Technical Analysis: Technical patterns, indicators with specific values. Note specialized indicators if present (e.g., ”AI Top/Bottom”, ”Bull Institution signals”).
@@ C.2 Writing Guidelines
  • Avoid phrases: ”according to”, ”information shows”, ”recent performance”, ”current situation”, ”comprehensive analysis”, ”buy”, ”sell”, ”hold”, ”clear position”, ”build position”, ”reduce position”, ”increase position”
  • Minimize transition words: ”but”, ”however”, ”then”, ”finally”
@@
Data (Total Score: 20) Extensive data (¿3 dimensions) 20
@@
Appendix D LLMs’ Response to users’ query
@@ D.1 FinSphere
    • Comprehensive data dimensions (¿3 dimensions)
@@ D.2 GPT-4o with few-shot
GPT-4o with few-shot’s analysis begins with vague conclusions: ”worth moderate attention in short term, should pay appropriate attention in medium to long term” lacks a clear stance; while data citations are complete, the analysis lacks depth; technical analysis is oversimplified and fails to fully interpret indicator meanings; the conclusion section lacks sufficient reasoning for investment recommendations. Specific scores are as follows:
@@
D.3 QWen2-72B with few-shot
@@
Appendix E A complete example of Stocksis
@@
Table 5: A complete example of Stocksis. Same sample as in Table 1. The average cost of comprehensive analysis written by experts is \$10 per query, with detailed cost breakdowns available in Appendix F.
@@
Appendix F Stocksis Collection and Evaluation Costs
To further quantify the value of Stocksis and AnalyScore, we disclose the expert-curated Stocksis dataset containing 5,000 entries and the human evaluation costs for assessing model outputs using AnalyScore. Excluding the costs of quantitative tool calls, the expert compilation of comprehensive analyses costs approximately \$75,000 (averaging \$15 per entry). The expert evaluation of 100 outputs from each of the models using AnalyScore criteria cost approximately \$12,000 (averaging \$10 per evaluation across all groups). These figures do not include the additional costs associated with expert development of the AnalyScore evaluation framework.
Appendix G Product Release Information
As FinSphere, a powerful stock analysis agent developed by a stock investment advisory company—we currently have a fully functional product demo and have made it freely available to the public in December 2024. Due to double-blind review requirements, we regrettably cannot showcase this promising product to the reviewers at this stage. However, we look forward to including access information for this free public tool in the final version of our paper.
Appendix H Detailed testing quiries
Here we disclose 100 queries used for testing and experts’ scores on FinSphere. For details, please check the Table 6, 7 and 8.
Appendix I Limitation
FinSphere’s performance depends on the accuracy and availability of real-time financial data, which may impact analysis reliability. The AnalyScore framework still requires human validation, limiting full automation. Additionally, FinSphere may struggle with nuanced financial reasoning and novel market events beyond its training. Future work should focus on improving real-time adaptability, reducing reliance on curated data, and expanding domain coverage for broader financial applications.